When 3.802 g of Na3PO4·12H2O react with excess BaCl2·2H2O,
how many moles of Ba3(PO4)2 would be produced?
1
Expert's answer
2020-05-12T11:35:16-0400

"2Na_3PO_4(aq)+3BaCl_2(aq) \rightarrow Ba_3(PO_4)_2(s)+6NaCl(aq)"


Convert grams of "Na_3PO_4 *12H_2O" to moles

"3.802 g (\frac{1 mole}{380.18 g})=0.01000 mol"


Convert moles of "Na_3PO_4*12H_2O" to moles of "Na_3PO_4"


"0.01000 mol Na_3PO_4*12H_2O (\frac{1 mole Na_3PO_4}{1 mole Na_3PO_4*12H_2O})=0.01000 mol Na_3PO_4"


Convert moles of "Na_3PO_4" to moles of "Ba_3(PO_4)_2"


"0.01000 mol Na_3PO_4(\frac{1 mole Ba_3(PO_4)_2}{2 mole Na_3PO_4})=0.005 mol Ba_3(PO_4)_2"


Convert moles of "Ba_3(PO_4)_2" to grams of "Ba_3(PO_4)_2"


"0.005 mol (\frac{601.92 g}{1 mole})=3.010 g Ba_3(PO_4)_2"